# Caught in the act: detections of recoiling supermassive black holes from simulations

### v1

@@ Abstract
We study the detectability of supermassive black holes (SMBHs) with masses of $M_{\bullet}\gtrsim 10^{9}\,\mathrm{M}_{\sun}$ displaced by gravitational wave recoil kicks $(v_{\rm kick}=0\mathrm{-}2000\,\mathrm{km}\,\mathrm{s}^{-1})$ in simulations of merging massive $(M_{\star}>10^{11}\,\mathrm{M}_{\sun})$ early-type galaxies. The used Ketju code combines the gadget-4 fast multiple gravity solver with accurate regularised integration and post-Newtonian corrections (up to PN3.5) around SMBHs. The ejected SMBHs carry clusters of bound stellar material (black hole recoil clusters, BRCs) with masses in the range of $10^{6}\lesssim M_{\text{BRC}}\lesssim 10^{7}\,\mathrm{M}_{\sun}$ and sizes of several $10\,\mathrm{pc}$. For recoil velocities up to 60% of the galaxy escape velocity, the BRCs are detectable in mock photometric images at a Euclid-like resolution up to redshift $z\sim 1.0$. By Monte Carlo sampling the observability for different recoil directions and magnitudes, we predict that in $\sim 20\%$ of instances the BRCs are photometrically detectable, most likely for kicks with SMBH apocenters less than the galaxy effective radius. BRCs occupy distinct regions in the stellar mass/velocity dispersion vs. size relations of known star clusters and galaxies. An enhanced velocity dispersion in excess of $\sigma\sim 600\,\mathrm{km}\,\mathrm{s}^{-1}$ coinciding with the SMBH position provides the best evidence for an SMBH-hosting stellar system, effectively distinguishing BRCs from other faint stellar systems. BRCs are promising candidates to observe the aftermath of the yet-undetected mergers of the most massive SMBHs and we estimate that up to 8000 BRCs might be observable below $z\lesssim 0.6$ with large-scale photometric surveys such as Euclid and upcoming high-resolution imaging and spectroscopy with the Extremely Large Telescope.
journal: ApJsoftware: Ketju (Mannerkoski et al., 2023; Rantala et al., 2017), gadget-4 (Springel et al., 2021), NumPy (Harris et al., 2020), SciPy (Virtanen et al., 2020), Matplotlib (Hunter, 2007), pygad (Röttgers et al., 2020), Stan (Stan Development Team, 2018), CmdStanPy (Stan Development Team, 2018), Arviz (Kumar et al., 2019), Synthesizer (Vijayan et al., 2020).
@@ 1 Introduction
Both observations (e.g. Tran et al., 2005; Trujillo et al., 2006; McIntosh et al., 2008; van Dokkum et al., 2010; van der Wel et al., 2014), and theoretical models (e.g. De Lucia et al., 2006; Naab et al., 2009; Johansson et al., 2012; Laporte et al., 2013; Behroozi et al., 2013; Rodriguez-Gomez et al., 2016; Moster et al., 2018, 2020) have long indicated that, in particular for massive early-type galaxies (ETGs) in dense environments, mergers of ETGs are frequent and dominate the mass assembly and size growth below redshifts of $z\lesssim 2$. Central supermassive black holes (SMBHs) are believed to reside at the centers of all massive galaxies (e.g. Kormendy & Ho, 2013), and if the host galaxies merge, subsequent sinking and coalescence of the SMBHs is expected to occur through a three-stage process (Begelman et al., 1980).
@@
Firstly, dynamical friction (Chandrasekhar, 1943) brings the SMBHs down from kiloparsec scales to parsec-scale separations. In the second phase, the SMBH separation is further reduced through sequential three-body interactions with the surrounding stellar component (Hills & Fullerton, 1980; Quinlan, 1996). In the final phase, at subparsec separations, gravitational wave (GW) emission becomes the dominant mechanism for the loss of the remaining orbital energy and angular momentum, thus driving the SMBHs to coalescence (Peters & Mathews, 1963). The sinking and resulting mergers of SMBHs provide plausible explanations for the observed low scatter in the SMBH scaling relations (e.g. Hirschmann et al., 2010; Jahnke & Macciò, 2011), the formation of the observed faint, diffuse central cores with missing starlight found at the centers of massive ETGs (e.g. Milosavljević & Merritt, 2001; Hoffman & Loeb, 2007; Rantala et al., 2018; Nasim et al., 2021), and repeated core scouring can also increase the effective radii of massive ETGs by up to a factor of two (e.g. Rantala et al., 2019, 2024b).
@@
Non-isotropic GW emission from the SMBH binary carries linear momentum away from the system (e.g. Bekenstein, 1973; González et al., 2007b), which at the time of coalescence, imparts a recoil (or kick) velocity to the remnant SMBH, varying from tens to thousands of kilometers per second in magnitude (Campanelli et al., 2007; González et al., 2007a). The magnitude of the recoil velocity depends on the progenitor SMBH mass difference and the relative difference between the SMBH spin vectors, with largest kicks occuring for equal-mass SMBHs with anti-aligned spins (González et al., 2007a). As the central escape velocity of a typical massive ETG is $v_{\rm esc}\sim 2000\,\mathrm{km}\,\mathrm{s}^{-1}$, a non-negligible fraction of galaxies might have ejected their SMBHs (e.g. Madau & Quataert, 2004; Mannerkoski et al., 2022), however not so many as to introduce considerable scatter in the observed SMBH scaling relations (Volonteri, 2007). Numerical simulations have also shown that the gravitational recoil-driven ejection of the coalesced SMBH can further increase the size of the galaxy core beyond that of the initially binary-scoured core (e.g. Gualandris & Merritt, 2008; Nasim et al., 2021; Khonji et al., 2024; Rawlings et al., 2025).
@@
Recent results from pulsar timing array (PTA) observations show tentative evidence for a stochastic gravitational wave background signal consistent with the merging of massive SMBHs $(M_{\bullet}\gtrsim 10^{8}M_{\odot})$ (Agazie et al., 2023; Xu et al., 2023; Zic et al., 2023; EPTA Collaboration et al., 2024).
@@
Presently, there is no direct observational confirmation for SMBH mergers (unlike stellar mass BHs) and resulting recoil kicks. However, several observational studies have hinted at the existence of recoiling SMBHs, generally through an offset from the center of the galaxy in either position, in velocity, or both (e.g. Komossa et al., 2008; Eracleous et al., 2012; Comerford & Greene, 2014; Pesce et al., 2021; Hogg et al., 2021; Chiaberge et al., 2025). Attributing an escaping SMBH to GW recoil, as opposed to a three-body interaction between a binary SMBH and a third SMBH, is not always straightforward, as evidenced in the recent detection by van Dokkum et al. (2023) of a runaway SMBH candidate identified by its interaction with the surrounding circumgalactic medium. Gravitationally recoiling SMBHs can also affect their accretion rates and thus quasar lifetimes, as the SMBHs are offset from the central gas density peak, potentially also inducing additional scatter in the observed SMBH scaling relations (e.g. Blecha & Loeb, 2008; Blecha et al., 2011).
@@
Using numerical simulations, Merritt et al. (2009) proposed that a recoiling SMBH would carry a retinue of bound stars with it (see also Boylan-Kolchin et al., 2004; O’Leary & Loeb, 2009). These stars would form a stellar system around the recoiling SMBH, which we will term a ‘black hole recoil cluster’ (BRC) in this work, with characteristic sizes and luminosities similar to globular clusters. However, in extreme cases, when kicked from massive galaxies, the BRC masses could even approach that of ultracompact dwarf galaxies. The defining characteristic of BRCs surrounding SMBHs and what sets them apart from other compact stellar systems is their very high internal stellar velocity dispersions, which scales with recoil velocity, and is typically two orders of magnitude higher than other compact stellar systems. Observational studies have already attempted to identify such massive clusters in massive cored ETGs (Burke-Spolaor et al., 2017; Gültekin et al., 2021). In addition, searches for lower-mass compact stellar systems have also been conducted in the Milky Way (e.g. Greene et al., 2021; Wu et al., 2024).
@@
In this paper, we utilise self-consistent gas-free simulations of merging ETGs with recoiling SMBHs that are run using the GADGET-4 based regularised tree code Ketju (Rantala et al., 2017, 2020; Mannerkoski et al., 2023). Specifically, we use the large simulation sample presented in Rawlings et al. (2025) in which the SMBH kick velocity was systematically varied to study its impact on the host galaxy. Earlier, Ketju has been used to study the formation of cores in gas-free simulations, but without the contribution of SMBH recoils (Rantala et al., 2018, 2019, 2024b; Mannerkoski et al., 2023; Partmann et al., 2024).
@@ 2.1 Simulation code
The simulations presented in this study were run with the new public version of the Ketju [^1] https://www.mv.helsinki.fi/home/phjohans/group-website/research/ketju/ code (Rantala et al., 2017, 2020; Mannerkoski et al., 2023), which is coupled to the GADGET-4 code (Springel et al., 2021). Ketju adds a small algorithmically regularised region (i.e., the Ketju region) around each SMBH, in which the dynamics of SMBH and the surrounding stellar component is integrated using the MSTAR integrator (Rantala et al., 2020). The dynamics of stellar particles beyond this small region and the dynamics of all dark matter particles are then computed with the GADGET-4 fast multipole method (FMM) gravity solver with second-order multipoles. We also employ hierarchical time integration, which allows for mutually symmetric interactions and manifest momentum conservation. Ketju also includes post-Newtonian (PN) correction terms up to order PN3.5 with optional spin terms between each pair of SMBHs (Blanchet, 2014) and the fitting formula of Zlochower & Lousto (2015) for recoil kick velocities.
@@ 2.2 Merger simulations
We use the gas-free isolated major merger simulations of massive ETGs from Rawlings et al. (2025) and summarise here the main properties of the simulation setup; for full simulation details, please refer to Rawlings et al. (2025).
Each progenitor galaxy is represented by an isotropic multicomponent system and consists of a stellar component with a total mass of $M_{\star}\sim 1.38\times 10^{11}\,\mathrm{M}_{\sun}$ embedded in a dark matter (DM) halo with a total mass of $M_{\mathrm{DM}}=2.5\times 10^{13}\,\mathrm{M}_{\sun}$ (Moster et al., 2010). At the center of the halo, we place an SMBH with a mass of $M_{\bullet,0}=2.93\times 10^{9}\,\mathrm{M}_{\sun}$ (Sahu et al., 2019) and zero initial velocity. Both the stellar and DM components each follow a Hernquist (1990) profile, with a scale radius of $a_{\star}=3.9\ \rm kpc$ and $a_{\star}=245\ \rm kpc$, respectively. The density profile $\rho_{i}$ of each component is then given by:
@@
The stellar and DM particle masses are $m_{\star}=5.0\times 10^{4}\,\mathrm{M}_{\sun}$ and $m_{\mathrm{DM}}=5.0\times 10^{6}\,\mathrm{M}_{\sun}$, respectively. The simulations are performed with Ketju, where SMBH-star and SMBH-SMBH interactions are non-softened, and all other interactions are softened: $\varepsilon_{\star}=2.5\,\mathrm{pc}$ for stellar particles, and $\varepsilon_{\mathrm{DM}}=200\,\mathrm{pc}$ for DM particles. All stellar particles beyond $3\varepsilon_{\star}=7.5\,\mathrm{pc}$ of an SMBH, and all DM particles, are integrated using gadget-4 (Springel et al., 2021).
@@
Whilst we neglect gas in these simulations, Ketju has previously been used to study SMBH merger timescales in hydrodynamical simulations, with the results from elliptical-elliptical mergers showing that the SMBH merger timescales and central stellar densities are similar for no-gas simulations and simulations that include both gas physics and models for stellar and BH feedback (Liao et al., 2024a, b). As a consequence, we do not expect gas physics to play a major role in the dynamics of the recoiling SMBHs in the systems we study.
@@ 2.3 Defining a black hole recoil cluster
We then require that the members of the BRC are strongly bound [^2] Note that to be bound, but not necessarily strongly, the particle binding energy must be greater than 0. to the SMBH, such that their binding energy $E_{\mathrm{b}}$ is greater than the ambient stellar velocity dispersion $\sigma_{\mathrm{amb}}$:
@@
Strongly-bound particles at the first SMBH apocentre and following pericentre are shown in Figure 1 as circle and square markers, respectively. There is negligible difference in $M_{\mathrm{BRC}}$ at these two moments, indicating that these strongly bound particles are travelling with the SMBH during its orbital excursion. At most, $M_{\mathrm{BRC}}\simeq 2\times 10^{-3}\,M_{\bullet}$. For recoils $v_{\mathrm{kick}}\gtrsim 540\,\mathrm{km}\,\mathrm{s}^{-1}$, $M_{\mathrm{BRC}}$ decreases with $v_{\mathrm{kick}}$, as the higher recoil velocity requires higher binding energies for the stellar particles to remain bound to the SMBH (see also Merritt et al., 2009). Conversely for $v_{\mathrm{kick}}\lesssim 540\,\mathrm{km}\,\mathrm{s}^{-1}$, $M_{\mathrm{BRC}}$ rises slightly. This occurs due to the relative velocity term $v_{i}$ in Equation 3: for low $v_{\mathrm{kick}}$, whilst there are many particles bound to the SMBH with $E_{\mathrm{b}}>0$, they do not necessarily have $E_{\mathrm{b}}>\sigma_{\mathrm{amb}}^{2}$, and are hence not strongly bound to the SMBH. As $v_{\mathrm{kick}}$ increases above $\sigma_{\mathrm{amb}}$, a greater fraction of stellar particles bound to the SMBH typically also have velocities above the ambient velocity dispersion. Until $v_{\mathrm{kick}}\simeq 540\,\mathrm{km}\,\mathrm{s}^{-1}\simeq 2\sigma_{\star,0}$ the recoil velocity of the SMBH is not so great so as to unbind a large amount of stellar mass from the recoiling SMBH (due to the kinetic energy of stellar particles increasing as $v_{i}^{2}$ from Equation 3), leading to the peak in $M_{\mathrm{BRC}}$ before declining for $v_{\mathrm{kick}}\gtrsim 2\sigma_{\star,0}$.
A large $M_{\mathrm{BRC}}$ is of interest when assessing the detectability of a recoiling SMBH with a non-zero orthogonal component of its velocity vector to the LOS (e.g. Merritt et al., 2009; O’Leary & Loeb, 2009) as the corresponding stellar luminosity may be detectable in photometric observations.

### v2

@@ Abstract
We study the detectability of supermassive black holes (SMBHs) with masses of $M_{\bullet}\gtrsim 10^{9}\,\mathrm{M}_{\sun}$ displaced by gravitational wave recoil kicks $(v_{\rm kick}=0\mathrm{-}2000\,\mathrm{km}\,\mathrm{s}^{-1})$ in simulations of merging massive $(M_{\star}>10^{11}\,\mathrm{M}_{\sun})$ early-type galaxies. The used Ketju code combines the gadget-4 fast multiple gravity solver with accurate regularised integration and post-Newtonian corrections (up to PN3.5) around SMBHs. The ejected SMBHs carry clusters of bound stellar material (black hole recoil clusters, BRCs) with masses in the range of $10^{6}\lesssim M_{\text{BRC}}\lesssim 10^{7}\,\mathrm{M}_{\sun}$ and sizes of several $10\,\mathrm{pc}$. For recoil velocities up to 60% of the galaxy escape velocity, the BRCs are detectable in mock photometric images at a Euclid-like resolution up to redshift $z\sim 1.0$. By Monte Carlo sampling the observability for different recoil directions and magnitudes, we predict that in $\sim 20\%$ of instances the BRCs are photometrically detectable, most likely for kicks with SMBH apocentres less than the galaxy effective radius. BRCs occupy distinct regions in the stellar mass/velocity dispersion vs. size relations of known star clusters and galaxies. An enhanced velocity dispersion in excess of $\sigma\sim 600\,\mathrm{km}\,\mathrm{s}^{-1}$ coinciding with the SMBH position provides the best evidence for an SMBH-hosting stellar system, effectively distinguishing BRCs from other faint stellar systems. BRCs are promising candidates to observe the aftermath of the yet-undetected mergers of the most massive SMBHs and we estimate that up to 8000 BRCs might be observable below $z\lesssim 0.6$ with large-scale photometric surveys such as Euclid and upcoming high-resolution imaging and spectroscopy with the Extremely Large Telescope.
journal: ApJsoftware: Ketju (mannerkoski2023; rantala2017), gadget-4 (springel2021), NumPy (harris2020), SciPy (virtanen2020), Matplotlib (hunter2007), pygad (rottgers2020), Stan (standevelopmentteam2018), CmdStanPy (standevelopmentteam2018), Arviz (kumar2019), Synthesizer (vijayan2021).
@@ 1 Introduction
Both observations (e.g. tran2005; trujillo2006; mcintosh2008; vandokkum2010; vanderwel2014), and theoretical models (e.g. delucia2006; Naab2009; Johansson2012; laporte2013; behroozi2013; Rodriguez-Gomez2016; moster2018; moster2020) have long indicated that, in particular for massive early-type galaxies (ETGs) in dense environments, mergers of ETGs are frequent and dominate the mass assembly and size growth below redshifts of $z\lesssim 2$. Central supermassive black holes (SMBHs) are believed to reside at the centers of all massive galaxies (e.g. kormendy2013), and if the host galaxies merge, subsequent sinking and coalescence of the SMBHs is expected to occur through a three-stage process (begelman1980).
@@
Firstly, dynamical friction (chandrasekhar1943) brings the SMBHs down from kiloparsec scales to parsec-scale separations. In the second phase, the SMBH separation is further reduced through sequential three-body interactions with the surrounding stellar component (hills1980; quinlan1996). In the final phase, at subparsec separations, gravitational wave (GW) emission becomes the dominant mechanism for the loss of the remaining orbital energy and angular momentum, thus driving the SMBHs to coalescence (peters1963). The sinking and resulting mergers of SMBHs provide plausible explanations for the observed low scatter in the SMBH scaling relations (e.g. hirschmann2010; jahnke2011), the formation of the observed faint, diffuse central cores with missing starlight found at the centers of massive ETGs (e.g. milosavljevic2001; hoffman2007; rantala2018; nasim2021b), and repeated core scouring can also increase the effective radii of massive ETGs by up to a factor of two (e.g. rantala2019; rantala2024).
@@
Non-isotropic GW emission from the SMBH binary carries linear momentum away from the system (e.g. bekenstein1973; gonzalez2007), which at the time of coalescence, imparts a recoil (or kick) velocity to the remnant SMBH, varying from tens to thousands of kilometers per second in magnitude (campanelli2007; gonzalez2007b). The magnitude of the recoil velocity depends on the progenitor SMBH mass difference and the relative difference between the SMBH spin vectors, with largest kicks occuring for equal-mass SMBHs with anti-aligned spins (gonzalez2007b). As the central escape velocity of a typical massive ETG is $v_{\rm esc}\sim 2000\,\mathrm{km}\,\mathrm{s}^{-1}$, a non-negligible fraction of galaxies might have ejected their SMBHs (e.g. madau2004; mannerkoski2022), however not so many as to introduce considerable scatter in the observed SMBH scaling relations (volonteri2007). Numerical simulations have also shown that the gravitational recoil-driven ejection of the coalesced SMBH can further increase the size of the galaxy core beyond that of the initially binary-scoured core (e.g. gualandris2008; nasim2021b; khonji2024; rawlings2025).
@@
Recent results from pulsar timing array (PTA) observations show tentative evidence for a stochastic gravitational wave background signal consistent with the merging of massive SMBHs $(M_{\bullet}\gtrsim 10^{8}M_{\odot})$ (agazie2023; 2023Xu; zic2023; antoniadis2024).
@@
Presently, there is no direct observational confirmation for SMBH mergers (unlike stellar mass BHs) and resulting recoil kicks. However, several observational studies have hinted at the existence of recoiling SMBHs, generally through an offset from the center of the galaxy in either position, in velocity, or both (e.g. komossa2008; Eracleous2012; comerford2014; pesce2021; hogg2021; Chiaberge2025). Attributing an escaping SMBH to GW recoil, as opposed to a three-body interaction between a binary SMBH and a third SMBH, is not always straightforward, as evidenced in the recent detection by vandokkum2023 of a runaway SMBH candidate identified by its interaction with the surrounding circumgalactic medium. Gravitationally recoiling SMBHs can also affect their accretion rates and thus quasar lifetimes, as the SMBHs are offset from the central gas density peak, potentially also inducing additional scatter in the observed SMBH scaling relations (e.g. blecha2008; blecha2011).
@@
Using numerical simulations, merritt2009 proposed that a recoiling SMBH would carry a retinue of bound stars with it (see also boylan2004; oleary2009). These stars would form a stellar system around the recoiling SMBH, which we will term a ‘black hole recoil cluster’ (BRC) in this work, with characteristic sizes and luminosities similar to globular clusters. However, in extreme cases, when kicked from massive galaxies, the BRC masses could even approach that of ultracompact dwarf galaxies. The defining characteristic of BRCs surrounding SMBHs and what sets them apart from other compact stellar systems is their very high internal stellar velocity dispersions, which scales with recoil velocity, and is typically two orders of magnitude higher than other compact stellar systems. Observational studies have already attempted to identify such massive clusters in massive cored ETGs (Burke-Spolaor2017; Gultekin2021). In addition, searches for lower-mass compact stellar systems have also been conducted in the Milky Way (e.g. Greene2021; Wu2024).
@@
In this paper, we utilise self-consistent gas-free simulations of merging ETGs with recoiling SMBHs that are run using the GADGET-4 based regularised tree code Ketju (rantala2017; rantala2020; mannerkoski2023). Specifically, we use the large simulation sample presented in rawlings2025 in which the SMBH kick velocity was systematically varied to study its impact on the host galaxy. Earlier, Ketju has been used to study the formation of cores in gas-free simulations, but without the contribution of SMBH recoils (rantala2018; rantala2019; rantala2024; mannerkoski2023; partmann2024).
@@ 2.1 Simulation code
The simulations presented in this study were run with the new public version of the Ketju [^1] https://www.mv.helsinki.fi/home/phjohans/group-website/research/ketju/ code (rantala2017; rantala2020; mannerkoski2023), which is coupled to the GADGET-4 code (springel2021). Ketju adds a small algorithmically regularised region (i.e., the Ketju region) around each SMBH, in which the dynamics of SMBH and the surrounding stellar component is integrated using the MSTAR integrator (rantala2020). The dynamics of stellar particles beyond this small region and the dynamics of all dark matter particles are then computed with the GADGET-4 fast multipole method (FMM) gravity solver with second-order multipoles. We also employ hierarchical time integration, which allows for mutually symmetric interactions and manifest momentum conservation. Ketju also includes post-Newtonian (PN) correction terms up to order PN3.5 with optional spin terms between each pair of SMBHs (blanchet2014) and the fitting formula of zlochower2015 for recoil kick velocities.
@@ 2.2 Merger simulations
We use the gas-free isolated major merger simulations of massive ETGs from rawlings2025 and summarise here the main properties of the simulation setup; for full simulation details, please refer to rawlings2025.
Each progenitor galaxy is represented by an isotropic multicomponent system and consists of a stellar component with a total mass of $M_{\star}\sim 1.38\times 10^{11}\,\mathrm{M}_{\sun}$ embedded in a dark matter (DM) halo with a total mass of $M_{\mathrm{DM}}=2.5\times 10^{13}\,\mathrm{M}_{\sun}$ (moster2010). At the center of the halo, we place an SMBH with a mass of $M_{\bullet,0}=2.93\times 10^{9}\,\mathrm{M}_{\sun}$ (sahu2019) and zero initial velocity. Both the stellar and DM components each follow a hernquist1990 profile, with a scale radius of $a_{\star}=3.9\ \rm kpc$ and $a_{\star}=245\ \rm kpc$, respectively. The density profile $\rho_{i}$ of each component is then given by:
@@
The stellar and DM particle masses are $m_{\star}=5.0\times 10^{4}\,\mathrm{M}_{\sun}$ and $m_{\mathrm{DM}}=5.0\times 10^{6}\,\mathrm{M}_{\sun}$, respectively. The simulations are performed with Ketju, where SMBH-star and SMBH-SMBH interactions are non-softened, and all other interactions are softened: $\varepsilon_{\star}=2.5\,\mathrm{pc}$ for stellar particles, and $\varepsilon_{\mathrm{DM}}=200\,\mathrm{pc}$ for DM particles. All stellar particles beyond $3\varepsilon_{\star}=7.5\,\mathrm{pc}$ of an SMBH, and all DM particles, are integrated using gadget-4 (springel2021).
@@
Whilst we neglect gas in these simulations, Ketju has previously been used to study SMBH merger timescales in hydrodynamical simulations, with the results from elliptical-elliptical mergers showing that the SMBH merger timescales and central stellar densities are similar for no-gas simulations and simulations that include both gas physics and models for stellar and BH feedback (liao2024a; liao2024b). As a consequence, we do not expect gas physics to play a major role in the dynamics of the recoiling SMBHs in the systems we study.
@@ 2.3 Defining a black hole recoil cluster
We then require that the members of the BRC are strongly-bound [^2] Note that to be bound, but not necessarily strongly, the particle binding energy must be greater than 0. to the SMBH, such that their binding energy $E_{\mathrm{b}}$ is greater than the ambient stellar velocity dispersion $\sigma_{\mathrm{amb}}$:
@@
Strongly-bound particles at the first SMBH apocentre and following pericentre are shown in Figure 1 as circle and square markers, respectively. There is negligible difference in $M_{\mathrm{BRC}}$ at these two moments, indicating that these strongly-bound particles are travelling with the SMBH during its orbital excursion. At most, $M_{\mathrm{BRC}}\simeq 2\times 10^{-3}\,M_{\bullet}$. For recoils $v_{\mathrm{kick}}\gtrsim 540\,\mathrm{km}\,\mathrm{s}^{-1}$, $M_{\mathrm{BRC}}$ decreases with $v_{\mathrm{kick}}$, as the higher recoil velocity requires higher binding energies for the stellar particles to remain bound to the SMBH (see also merritt2009). Conversely for $v_{\mathrm{kick}}\lesssim 540\,\mathrm{km}\,\mathrm{s}^{-1}$, $M_{\mathrm{BRC}}$ rises slightly. This occurs due to the relative velocity term $v_{i}$ in Equation 3: for low $v_{\mathrm{kick}}$, whilst there are many particles bound to the SMBH with $E_{\mathrm{b}}>0$, they do not necessarily have $E_{\mathrm{b}}>\sigma_{\mathrm{amb}}^{2}$, and are hence not strongly-bound to the SMBH. As $v_{\mathrm{kick}}$ increases above $\sigma_{\mathrm{amb}}$, a greater fraction of stellar particles bound to the SMBH typically also have velocities above the ambient velocity dispersion. Until $v_{\mathrm{kick}}\simeq 540\,\mathrm{km}\,\mathrm{s}^{-1}\simeq 2\sigma_{\star,0}$ the recoil velocity of the SMBH is not so great so as to unbind a large amount of stellar mass from the recoiling SMBH (due to the kinetic energy of stellar particles increasing as $v_{i}^{2}$ from Equation 3), leading to the peak in $M_{\mathrm{BRC}}$ before declining for $v_{\mathrm{kick}}\gtrsim 2\sigma_{\star,0}$.
A large $M_{\mathrm{BRC}}$ is of interest when assessing the detectability of a recoiling SMBH with a non-zero orthogonal component of its velocity vector to the LOS (e.g. merritt2009; oleary2009) as the corresponding stellar luminosity may be detectable in photometric observations.